OpenStudy (anonymous):

Use the triangle at the right. Find the length of the missing side. 1. a = 16, b = 63 2. b = 2.1, c = 2.9

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

pythagoras \[a^2+b^2=c^2\]

OpenStudy (anonymous):

\[16^2+63^2=c^2\] \[c=\sqrt{16^2+63^2}\]

OpenStudy (anonymous):

i just need help with number 2